OpenStudy (anonymous):

Find the standard deviation. Round to one more place than the data. 2, 3, 14, 20, 18, 20, 19, 8, 12

OpenStudy (loser66):

7.131698

OpenStudy (anonymous):

The formula for standard deviation is \[\sqrt{\frac{ \sum }{ n }}\] Where \[\sum\] is the sum of all the values given and "n" is the number of values. Just add the values together and divide them by the number of values you have and then take the square root of it.